OpenStudy (anonymous):

Simplify the following expression: (x − 8y) - (2x − 4y). If the final answer is written in the form Ax + By, what is the value of B?

OpenStudy (anonymous):

(x-8y)-(2x-4y) = x-2x -8y+4y = -x -4y. So B would equal -4.